bu_cms_history/oSLB_July_2007_Test_Summary

SiteMap (Historical BU CMS wiki main page)

We have completed our oSLB testing. We have fibers installed between HO crate 7 and RPC as in the diagram:

  https://cmsdaq.cern.ch/elog/HCAL/070703_150038/diagram_HCAL2RPC.pdf

Summary of test results:

Other links have problems which we will investigate

With help from Andrei Holzner, we cabled a BC0 input to the RPC and HO TTCci interfaces. Using the TTCci configuration for hcal:

  http://joule.bu.edu/~hazen/CMS/oSLB_Jul07/eric_config2.txt

we send BC0 with no delay, and the test pattern after 50 clocks. For the short (10-12m) fibers, the RPC sees BC0 at 6-7 clocks after the RPC BC0, and the muon hits at 66-67 clocks.

The 6 clock BC0 delay includes some HTR processing time, 1-2 clocks in the oSLB and 2-3 clocks for the GOL+TLK deserializer plus fiber delay. The data comes about 10 clocks later, which is some combination of the HTR input pattern ram latency and the TPG path.

Work to be done:

o Understand non-working links (I'll try to do this today or tomorrow) o Define xDAQ control for oSLB in detail o Update oSLB firmware to support a new data format to be detailed by Karol

It is likely we will re-convene for future testing near the Sept. global run.